The sum of two numbers is 71. The second is 6 more than 4 times the first. What are the two numbers?
thanks for the help.

Answer by tutorcecilia(2152)

x+y=71
x+(4x+6)=71
5x+6=71 [solve for x]
5x=71-6
5x=65
x=13
.
Therefore,
x+y=71
13+y=71
y=71-13
y=58
or y=4x+6=4(13)+6=58
or 13+58=71